OpenStudy (anonymous):

from a standard deck, what is the probability of picking either an eighy or a queen

OpenStudy (cwrw238):

there are 4 eight's and 4 queens in a pack. Probability(8) = 4/52 or 1/13 Probability(Q) = 1/13

OpenStudy (cwrw238):

can you proceed from here? what do you do to the 2 probabilities to get P(8 or Q)?